OpenStudy (calculusxy):

If x=y-7, what is the value of 3x-3 in terms of y?

OpenStudy (anonymous):

hi x=y-7 so 3x=3y-21 so 3x-3=3y-24

OpenStudy (calculusxy):

Why is it 3y?

OpenStudy (anonymous):

3x=3(y-7)=3y-21